OpenStudy (anonymous):

Simplify 2√5u + 6√45u

OpenStudy (anonymous):

take out a common factor. first simplify 5sqrt(45); what can that be simplified to?

OpenStudy (anonymous):

45 would go to 9*5

OpenStudy (anonymous):

Ok, lets check it out.

OpenStudy (anonymous):

\[2 \sqrt(5u)+6 \sqrt(45u) = 2 \sqrt(5)\sqrt(u) + 6 \sqrt(9) \sqrt(5) \sqrt(u)\]

OpenStudy (anonymous):

\[2 \sqrt(5u)+6 \sqrt(45u) = 2 \sqrt(5)\sqrt(u) + 6 \sqrt(9) \sqrt(5) \sqrt(u)= 2 \sqrt(5) \sqrt(u) +18 \sqrt(5) \sqrt(u) = \sqrt(u)(20 \sqrt(5))\]

OpenStudy (anonymous):

= \[\sqrt(u) (20 \sqrt(5))\]